# Double-Estimation-Friendly Inference for High Dimensional Misspecified Measurement Error Models

### v1

@@ 3 Sparsity-Adaptive Procedure
In high dimensional settings, the statistic proposed above is valid when Condition 5 holds, which requires that the coefficients $\bm{\gamma}_{0}$ and $\bm{\theta}_{0}$ are both sparse enough. However, this assumption might not hold in practice. For instance, biologists have found examples from genome-wide gene expression profiling that all genes are believed to affect a common disease marker, thereby violating the sparsity assumption. Thus, we need to consider a methodology that can work when the sparsity assumption is violated, specifically when either $\bm{\gamma}_{0}$ or $\bm{\theta}_{0}$ is dense. Zhu and Bradic (2017), Zhu and Bradic (2018a), and Zhu and Bradic (2018b) have proposed methods for statistical inference in non-sparse high dimensional linear models. Bradic et al. (2020) have generalized these methods to linear mixed models. Instead of using the penalty-based estimator directly for inference, they modified the Dantzig selector estimator proposed by Candes et al. (2007) and made it adaptive for inference in high dimensional models. We leverage their estimation procedure and generalize it to models with measurement error.
@@
Remark. Condition 10 is also named as restricted eigenvalue condition. When the covaraites are sub-Gaussian distributed, Rudelson and Zhou (2012) show that the restricted eigenvalue condition holds with probability tending to 1 when sample size $n$ goes to infinity.
@@
Suppose that 8 - 11 hold. Then under the null hypothesis, it holds that $T_{DF}\stackrel{{\scriptstyle d}}{{\longrightarrow}}N(0,1)$ if model (c) or (d) holds.
@@ 4.2 Performance under High Dimensional Setting
We then consider high dimensional setting in the following examples. In the following example, sample size $n$ is set to be $200$. In Example 1 and Example 2, We show the performance of $T_{DF}$ under model misspecification with the parameters still being sparse. In these examples, we use the Lasso estimator. We use $T_{DFLA}$ to represent our statistics with the lasso estimator, In Example 3, The performance of $T_{DF}$ using the sparsity-adaptive estimator is shown, under models with dense parameters. Use $T_{DFDE}$ to represent the statistics with the sparsity-adaptive estimator. We compare the performance of the newly proposed test statistics with the corrected decorrelated score test (CDST) statistics proposed by Li et al. (2021).
@@
Simulation results are shown in left penal in Figure 2. CDST performs similar to our method under this setting.
@@
Model 2a. $Y_{i}=X_{i}\beta+\widetilde{Z}_{i}^{\top}\bm{\gamma}+\varepsilon_{i},\widetilde{Z}_{i,j}=\dfrac{exp(Z_{i,j})-1}{exp(Z_{i,j})+1}.$ $\bm{\gamma}_{1},\cdots,\bm{\gamma}_{10}$ are from Uniform(0,1).
@@ 6 Discussion
7 Appendix
7.1 Proof of Theorem 1
@@ 7.3 Proof of Theorem 3
Proof of Theorem 3: Firstly, assume that model (2) holds while model (LABEL:5.220220501) is misspecified. In this situation, we still have
@@
(i) follows from Lemma 3. Next we consider the situation that model (LABEL:5.220220501) holds while model (2) is misspecified. In this situation, we have:
@@
7.4 Proof of Theorem 4
Assume that model (2) holds while model (LABEL:5.220220501) is misspecified. To look at local power of the statistics, denote $\Delta\beta=\beta_{0}-\beta^{*}=n^{-1/2}c$. In this situation, we have

### v3

@@ 3 Sparsity-Adaptive Procedure
In high dimensional settings, the statistic proposed above is valid when Condition (C5) holds, which requires that the coefficients $\bm{\gamma}_{0}$ and $\bm{\theta}_{0}$ are both sparse enough. However, this assumption might not hold in practice. For instance, biologists have found examples from genome-wide gene expression profiling that all genes are believed to affect a common disease marker, thereby violating the sparsity assumption. Thus, we need to consider a methodology that can work when the sparsity assumption is violated, specifically when either $\bm{\gamma}_{0}$ or $\bm{\theta}_{0}$ is dense. Zhu and Bradic (2017), Zhu and Bradic (2018a), and Zhu and Bradic (2018b) have proposed methods for statistical inference in non-sparse high dimensional linear models. Bradic et al. (2020) have generalized these methods to linear mixed models. Instead of using the penalty-based estimator directly for inference, they modified the Dantzig selector estimator proposed by Candes et al. (2007) and made it adaptive for inference in high dimensional models. We leverage their estimation procedure and generalize it to models with measurement error.
@@
Remark. Condition (C10) is also named as restricted eigenvalue condition. When the covaraites are sub-Gaussian distributed, Rudelson and Zhou (2012) show that the restricted eigenvalue condition holds with probability tending to 1 when sample size $n$ goes to infinity.
@@
Suppose that (C8) - (C11) hold. Then under the null hypothesis, it holds that $T_{DF}\stackrel{{\scriptstyle d}}{{\longrightarrow}}N(0,1)$ if model (c) or (d) holds.
@@ 4.2 Performance under High Dimensional Setting
We then consider high dimensional setting in the following examples. In the following examples, sample size $n$ is set to $200$. In Example 1 and Example 2, we show the performance of $T_{DF}$ under model misspecification with the parameters still being sparse. In these examples, we use the Lasso estimator. We use $T_{DFLA}$ to represent our statistics with the lasso estimator. In Example 3, the performance of $T_{DF}$ using the sparsity-adaptive estimator is shown, under models with dense parameters. Use $T_{DFDE}$ to represent the statistics with the sparsity-adaptive estimator. We compare the performance of the newly proposed test statistics with the corrected decorrelated score test (CDST) statistics proposed by Li et al. (2021).
@@
Simulation results are shown in the left penal in Figure 2. CDST performs similar to our method under this setting.
@@
Model 2a. $Y_{i}=X_{i}\beta+\widetilde{Z}_{i}^{\top}\bm{\gamma}+\varepsilon_{i},\widetilde{Z}_{i,j}=\dfrac{e^{Z_{i,j}}-1}{e^{Z_{i,j}}+1}.$ $\bm{\gamma}_{1},\cdots,\bm{\gamma}_{10}$ are from Uniform(0,1).
@@ 6 Discussion
Appendix
Proof of Theorem 1
@@
Proof of Theorem 3: Firstly, assume that $Y$- model in (2) holds while $X$- model in (2) is misspecified. In this situation, we still have
@@
(i) follows from Lemma 3. Next we consider the situation that $X$- model in (2) holds while $Y$- model in (2) is misspecified. In this situation, we have:
@@
Proof of Theorem 4
Assume that $Y$- model in (2) holds while model $X$-model in (2) is misspecified. To look at local power of the statistics, denote $\Delta\beta=\beta_{0}-\beta^{*}=n^{-1/2}c$. In this situation, we have